OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a red marble, replacing it, and then picking a black marble?

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

how many marbles total?

OpenStudy (anonymous):

15

OpenStudy (misty1212):

ok and haow many are red?

OpenStudy (anonymous):

5

OpenStudy (misty1212):

then what is the probability of picking a red one first?

OpenStudy (anonymous):

1/3 simplified

OpenStudy (misty1212):

right then you put the marble back, so there are still 15, how many are black?

OpenStudy (anonymous):

4

OpenStudy (misty1212):

so what is the probability of picking a black one ?

OpenStudy (anonymous):

4/15?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

to get your answer, multiply

OpenStudy (misty1212):

\[\frac{1}{3}\times \frac{4}{15}\]

OpenStudy (anonymous):

oh. i didnt know to multiply that. thank you!!